OpenStudy (anonymous):

factoring i already factored but do not know what to do next i got 3x(n^4-7x^2) please help!

OpenStudy (precal):

what is the original problem?

OpenStudy (anonymous):

3xn^4 - 27x^3

OpenStudy (precal):

\[3n^4-27n^3\] is this your problem?

OpenStudy (anonymous):

no 3xn to the 4

OpenStudy (dumbcow):

the inside is a difference of squares however it should be 3x(n^4 -9x^2) .....9*3 = 27

OpenStudy (precal):

\[3xn^4-27n^3\] this is it?

OpenStudy (precal):

I think I will let dumbcow solve this one.....

OpenStudy (dumbcow):

a^2 - b^2 = (a+b)(a-b) --> n^4-9x^2 = (n^2 +3x)(n^2 -3x)